# Wireless Josephson parametric amplifier above 20 GHz

### v1

@@ I Introduction
Conventional superconducting qubit platforms for large-scale quantum computation Google Quantum AI and Collaborators (2025); Gao et al. (2025) employ sub-10 GHz qubits operating below 20 mK in dilution refrigerators. At these temperatures, limited cooling power presents a significant challenge to scalability. Raising the operating temperature to the range of 100 mK–1 K offers substantially greater cooling headroom, enabling the operation of more qubits. However, this increase in temperature also elevates the excited-state population of sub-10 GHz qubits, necessitating more complex state preparation protocols such as autonomous coolingMagnard et al. (2018) or measurement-based resetJohnson et al. (2012). Additionally, thermal population in sub-10 GHz readout resonators contributes to qubit dephasing via photon shot noiseKrantz et al. (2019).
@@
To mitigate these effects, circuit QED systems operating at elevated temperatures must also transition to higher frequencies—typically in the 20–100 GHz range. Recent studies have demonstrated qubits operating above 20 GHz using advanced Josephson junctions and weak-links with superconducting Nb Anferov et al. (2024a, 2025) and TiN Purmessur et al. (2025). In parallel, readout resonators are also being pushed to higher frequencies to suppress thermal photon-induced dephasing. High-frequency resonators coupled to conventional frequency qubits are also under active investigation, as large detuning between qubit and readout frequencies can reduce measurement-induced state transitions, thereby enhancing readout fidelity and repeatability Swiadek et al. (2024); Nesterov and Pechenezhskiy (2024); Kurilovich et al. (2025). These developments collectively call for the development of ultra-low-noise amplification systems operating above 20 GHz to enable fast, high-fidelity qubit readout.
@@
State-of-the-art superconducting qubit measurement chains typically employ Josephson parametric amplifiers (JPAs) as the first stage of nearly quantum-limited amplification Aumentado (2020); Roy and Devoret (2016). However, conventional JPA designs operating below 12 GHz face significant challenges when scaled to higher frequencies. Standard packaging techniques—such as wirebonding JPA chips to printed circuit boards (PCBs)—introduce loss and impedance mismatches that worsen at high frequencies. Moreover, parasitic modes in these packages can interfere with the JPA mode and degrade noise performance. While some of these limitations can be mitigated Anferov et al. (2024a), scaling beyond 20 GHz is more readily achieved with designs that eliminate wirebonds and lossy PCB materials altogether. A promising precedent was established by Narla et al. Narla et al. (2014), who demonstrated a wireless JPA coupling scheme using a rectangular waveguide to realize parametric amplification at 10.2 GHz and below. In this work, we show that such wireless coupling schemes can be extended to realize parametric amplification at higher frequencies required for next-generation circuit QED systems.
@@
We report the design and characterization of a wireless Josephson parametric amplifier (WJPA) that operates in the K-band (18–26 GHz). The device, realized with aluminum/aluminum-oxide JJs, achieved more than 20 dB of gain over a tunable frequency range of 21 to 23.5 GHz, with an average dynamic bandwidth of 3 MHz. The WJPA, operated as a single pump four-wave mixing amplifier, achieved a 1 dB input compression power at 20 dB gain of dBm. Using Y-factor and qubit-based photon number calibration techniques, the system-added noise with the WJPA was measured to be approximately two photons at 22 GHz.
II WJPA device design and fabrication
@@
The WJPA device is shown in Fig. 1(a-d), with equivalent circuit diagram depicted in Fig. 1(e). The device is realized by a planar on-chip circuit mounted in a 3D copper package, with microwave signals applied through a rectangular waveguide launch. The rectangular waveguide dimensions were chosen to be 10.67 mm by 4.32 mm, corresponding to the WR42 standard rectangular waveguide specification, for frequency range of 18 to 26 GHz (K-band). The package is designed to be attached to a commercial WR42-2.92mm coaxial cable adapter (not shown) which is then connected to other microwave components. Fig. 1(b) shows a top view of the chip sitting in the copper package where the left opening mates with the WR42 adapter. Inspired by previous work on millimeter-wave superconducting circuits Anferov et al. (2024b, 2025), this design avoids wirebonds by using a lithographically defined transition from the waveguide to a slotline transmission line. The transition takes the form from Anferov et al. Anferov et al. (2025), where is the smaller rectangular waveguide dimension (4.32 mm for WR42), is the slotline gap width, and mm is the transition length in our optimized design. With electromagnetic simulation in Ansys HFSS, the taper was found to have a return loss below 10 dB and insertion loss under 1 dB over the target frequency band of 18 to 26 GHz. The slotline, defined on a sapphire substrate with superconducting ground planes, was designed with a gap width = 200 m, yielding a characteristic impedance of approximately . The slotline is shorted at one end, and a tunable lumped-element resonator is capacitively coupled at a distance mm, corresponding to a quarter-wavelength at frequency of approximately GHz.
@@
An optical micrograph of the lumped-element resonator are shown in Fig. 1(c). The resonator is composed of an Al/AlOx Josephson junction (JJ) SQUID and interdigitated finger capacitors . The lithographically defined coupling capacitors were designed to achieve a coupling quality factor Q 100, and corresponding coupling rate 200 MHz. In electromagnetic simulations, we found variation of when varying the Josephson inductance from 100 to 200 pH. Crucially, we found that the junction capacitance , often ignored in lower frequency applications, must be taken into account for higher-frequency designs. This is because becomes comparable to as the targeted resonant frequency approaches the JJ plasma frequency, GHz.
@@
Additionally, the junction participation ratio is a crucial parameter for ensuring that this nonlinear resonator can operate reliably as a parametric amplifier for amplifying signals from qubits. Here corresponds to the effective inductance of the resonant mode, which includes the contribution of the non-negligible parasitic inductance of the shunt capacitors. As described in supplementary Sec. I, we estimate from black-box quantization analysis of the simulated admittance Manucharyan et al. (2007); Nigg et al. (2012); Frattini et al. (2018). This gives a quality factor-participation ratio product , which ensures, according to previous results, that the device should reliably achieve gain dB Manucharyan et al. (2007); Narla et al. (2014); Schackert (2013).
@@
The WJPA chip is fabricated by first defining the slotline in tantalum on a sapphire wafer using photolithography and wet etching, followed by the fabrication of aluminum/aluminum-oxide Josephson junctions with the Dolan bridge technique. The effective Josephson inductance of the SQUID loop (Fig. 1(d)) is estimated to be about 90 pH at zero flux. During packaging, indium is applied around the chip edges to provide mechanical stability and suppress parasitic package modes.
III WJPA linear response and gain characterization
@@
To characterize the amplifier’s dynamic range, we measure gain compression as a function of input probe power for various fixed pump powers, as shown in Fig. 2(d). The 1 dB compression points (), marked with circles, is about -125 dBm at 20 dB of maximum gain. increases as the gain is decreased, and the extracted values decrease approximately linearly with pump power on a logarithmic scale, with a fitted slope of 0.6. This trend aligns with theoretical analysis and prior experimental observations where gain saturation arises from Kerr-induced detuning and pump depletion Eichler and Wallraff (2014); Planat et al. (2019).
@@
IV WJPA noise characterization
Characterization of the added noise of a Josephson amplifier is crucial for its application in qubit readout. In this work, we implemented two methods to characterize the noise performance of the WJPA and cross checked the results. The experiment wiring configurations for both methods can be found in supplementary Sec. II. We start by using the Y-factor technique with a variable temperature source (VTS) Malnou et al. (2024), consisting of a 50 matched microwave attenuator thermally anchored to a copper bracket with a heater and ruthenium oxide thermometer, and thermally isolated from the mixing chamber stage of the dilution refrigerator. By varying the VTS temperature, thermal noise is generated corresponding to a known photon occupation number , where , are Planck and Boltzmann constants, is the frequency and is the attenuator temperature.
@@ IV WJPA noise characterization
Fig. 3(a) shows the measured output noise power versus frequency for various . Unlike the expectation for an ideal linear amplifier where would increase with , we observe a reduction in output noise with increasing temperatures. As discussed in Ref. Malnou et al., 2024, this behavior arises because of gain compression, from the broadband Johnson noise at elevated saturating the amplifier and reducing its gain.
@@
This gain compression is visualized in Fig. 3(b), where we plot the gain versus frequency measured by the VNA for each . The gain decreases with increasing , dropping from 21.6 dB at 0.1 K to 12.9 dB at 1.75 K. To correctly extract the added noise, we account for this gain reduction by scaling the noise spectrum with the measured gain following the procedure discussed in Ref. Malnou et al., 2024, and further described in supplementary Sec. III. The resulting system-added noise as a function of frequency is shown in Fig. 3(c). It shows that across the amplifier bandwidth, the system added noise is approximately 2 photon.
@@
Next, we implemented a qubit-based noise calibration approach using a transmon-cavity system Macklin et al. (2015); Bultink et al. (2018) with a readout cavity frequency of GHz. Details of this method is described in supplementary Sec. IV. Fig. 3(d) shows the calibrated noise spectrum referenced to the input of the WJPA, acquired at 22 GHz. When the WJPA is turned off, the system noise is dominated by the HEMT, measured to be approximately 10.3 photons, or 10.9 K at 22 GHz. With the WJPA turned on, the calibrated added noise drops to a system noise of photons, or K. We can also express these results in terms of a measurement efficiency of with the WJPA turned on, compared to 9.7% with the WJPA turned off.
@@
To summarize, both noise calibration methods indicate the system added noise with the WJPA operating of approximately 2 photon. This value is likely limited by losses in coaxial cables and circulators between the VTS/qubit measurement plane and the WJPA port. Overall, this noise performance confirms the WJPA’s capability to operate near the quantum limit and substantially improve fidelity of measurement in high-frequency readout applications.
V Conclusion
@@
In summary, we have demonstrated a wireless, waveguide-coupled Josephson parametric amplifier (WJPA) that features a lumped-element resonator embedded with a JJ SQUID and coupled to the input/output port via a lithographically defined waveguide-to-slotline coupling structure. Operated in four-wave mixing mode with a single pump tone, the WJPA achieves over 20 dB of gain across a tunable range from 21 GHz to 23.5 GHz, with typical 3 MHz bandwidth. Using both Y-factor measurements and a qubit-based noise calibration, we extract an system added noise of approximately two photons. These results represent the first demonstration of a Josephson parametric amplifier operating above 20 GHz.
While this work realized a WJPA with Al/AlOx junctions, the design is readily implementable with Josephson elements realized with superconductors such as niobium, that can operate at high temperature KAnferov et al. (2025). Thus, this amplifier design is suitable for high-fidelity, fast qubit readout in high-temperature superconducting qubit systems.
Acknowledgment

### v3

@@ Abstract
Conventional superconducting qubit platforms for large-scale quantum computation google2025; Gao2025establishing employ sub-10 GHz qubits operating below 20 mK in dilution refrigerators. At these temperatures, limited cooling power presents a significant challenge to scalability. Raising the operating temperature to the range of 100 mK–1 K offers substantially greater cooling headroom, enabling the operation of more qubits. However, this increase in temperature also elevates the excited-state population of sub-10 GHz qubits, necessitating more complex state preparation protocols such as autonomous coolingMagnard2018fast or measurement-based resetJohnson2012herald. Additionally, thermal population in sub-10 GHz readout resonators contributes to qubit dephasing via photon shot noisekrantz2019quantum.
@@
To mitigate these effects, circuit QED systems operating at elevated temperatures must also transition to higher frequencies—typically in the 20–100 GHz range. Recent studies have demonstrated qubits operating above 20 GHz using advanced Josephson junctions and weak-links with superconducting Nb anferov2024super; anferov2024millimeter and TiN purmessur2025operation. In parallel, readout resonators are also being pushed to higher frequencies to suppress thermal photon-induced dephasinganferov2024super. High-frequency resonators coupled to conventional frequency qubits are also under active investigation, as large detuning between qubit and readout frequencies can reduce measurement-induced state transitions, thereby enhancing readout fidelity and repeatability Swiadek2024enhancing; Nesterov2024measurement; kurilovich2025high. These developments collectively call for the development of ultra-low-noise amplification systems operating above 20 GHz to enable fast, high-fidelity qubit readout.
@@
Kinetic-inductance traveling-wave parametric amplifiers (KITWPAs) have demonstrated amplification up to 34 GHz for astronomy applicationsshu2021nonlinearity; Tan2024operation. However, their added noise was not characterized and they are not in widespread use for superconducting qubit measurement. Instead, state-of-the-art superconducting qubit measurement systems typically employ Josephson parametric amplifiers (JPAs) as the first stage of nearly quantum-limited amplification aumentado2020; Roy2016. However, conventional JPA designs operating below 12 GHz face significant challenges when scaled to higher frequencies. Standard packaging techniques—such as wirebonding JPA chips to printed circuit boards (PCBs)—introduce loss and impedance mismatches that worsen at high frequencies. Moreover, parasitic modes in these packages can interfere with the JPA mode and degrade noise performance. While some of these limitations can be mitigated anferov2024super, scaling beyond 20 GHz is more readily achieved with designs that eliminate wirebonds and lossy PCB materials altogether. A promising precedent was established by Narla et al. narla2014, who demonstrated a wireless JPA coupling scheme using a rectangular waveguide to realize parametric amplification at 10.2 GHz and below. In this work, we show that such wireless coupling schemes can be extended to realize parametric amplification at higher frequencies required for next-generation circuit QED systems.
@@
We report the design and characterization of a wireless Josephson parametric amplifier (WJPA) that operates in the K-band (18–26 GHz). The device, realized with aluminum/aluminum-oxide JJs, achieved more than 20 dB of gain over a tunable frequency range of 21 to 23.5 GHz, with an average dynamic bandwidth of 3 MHz. The WJPA, operated as a single pump four-wave mixing amplifier, achieved a 1 dB input compression power at 20 dB gain of dBm. Using Y-factor and qubit-based photon number calibration techniques, the system added noise with the WJPA was measured to be approximately two photons at 22 GHz.
@@
The WJPA device is shown in Fig. 1(a-d), with equivalent circuit diagram depicted in Fig. 1(e). The device is realized by a planar on-chip circuit mounted in a 3D copper package, with microwave signals applied through a rectangular waveguide launch. The rectangular waveguide dimensions were chosen to be 10.67 mm by 4.32 mm, corresponding to the WR42 standard rectangular waveguide specification, for frequency range of 18 to 26 GHz (K-band). The package is designed to be attached to a commercial WR42-2.92mm coaxial cable adapter (not shown) which is then connected to other microwave components. Fig. 1(b) shows a top view of the chip sitting in the copper package where the left opening mates with the WR42 adapter. Inspired by previous work on millimeter-wave superconducting circuits anferov2024low; anferov2024millimeter, this design avoids wirebonds by using a lithographically defined transition from the waveguide to a slotline transmission line. The transition takes the form from Anferov et al. anferov2024low, where is the smaller rectangular waveguide dimension (4.32 mm for WR42), is the slotline gap width, and mm is the transition length in our optimized design. An illustration of these geometric dimensions is depicted in the Supplementary Fig. S1(a). With electromagnetic simulation in Ansys HFSS, the taper was found to have a return loss below 10 dB and insertion loss under 1 dB over the target frequency band of 18 to 26 GHz. The slotline, defined on a sapphire substrate with superconducting ground planes, was designed with a gap width m, yielding a characteristic impedance of approximately . The slotline is shorted at one end, and a tunable lumped-element resonator is capacitively coupled at a distance mm, corresponding to a quarter-wavelength at frequency of approximately GHz.
@@
An optical micrograph of the lumped-element resonator are shown in Fig. 1(c). The resonator is composed of an Al/AlOx Josephson junction (JJ) SQUID and interdigitated finger capacitors . The lithographically defined coupling capacitors were designed to achieve a coupling quality factor 100, and corresponding coupling rate MHz. In electromagnetic simulations, we found variation of when varying the Josephson inductance from 90 to 200 pH. Crucially, we found that the junction capacitance , often ignored in lower frequency applications, must be taken into account for higher-frequency designs. This is because , which ranges between 290 to 130 fF for the range above, becomes comparable to fF as the targeted resonant frequency approaches the JJ plasma frequency, GHz.
@@
Additionally, the junction participation ratio is a crucial parameter for ensuring that this nonlinear resonator can operate reliably as a parametric amplifier for amplifying signals from qubits. Here corresponds to the effective inductance of the resonant mode, which includes the contribution of the non-negligible parasitic inductance ( pH) of the shunt capacitors. As described in supplementary Sec. I, we estimate from black-box quantization analysis of the simulated admittance Manucharyan2007; Nigg2012; Frattini2018. This gives a quality factor-participation ratio product , which ensures, according to previous results, that the device should reliably achieve gain dB Manucharyan2007; narla2014; schackert2013practical.
@@
The WJPA chip is fabricated by first defining the slotline in tantalum (though other superconducting materials could also be used) on a sapphire wafer using photolithography and wet etching, followed by the fabrication of aluminum/aluminum-oxide Josephson junctions with the Dolan bridge technique. This SQUID loop (Fig. 1(d)) yields a total critical current of 3.7 A, estimated from the room temperature resistance, corresponding to an effective Josephson inductance of approximately 90 pH at zero flux. During packaging, indium is applied around the chip edges to provide mechanical stability and suppress parasitic package modes.
@@
To characterize the amplifier’s dynamic range, we measure gain compression as a function of input probe power for various fixed pump powers, as shown in Fig. 2(d). The 1 dB compression points (), marked with circles, is about -125 dBm at 20 dB of maximum gain. increases as the gain is decreased, and the extracted values decrease approximately linearly with pump power on a logarithmic scale, with a fitted slope of . This trend aligns with theoretical analysis and prior experimental observations where gain saturation arises from Kerr-induced detuning and pump depletion Eichler2014; Planat2019understand. To further enhance the amplifier’s dynamic range and reduce pump-induced dephasing in qubit readout applications, a double pump configuration narla2014 can be employed, which is compatible with the present design.
@@
Characterization of the added noise of a Josephson amplifier is crucial for its application in qubit readout. In this work, we implemented two methods to characterize the noise performance of the WJPA and cross checked the results. The experiment wiring configurations for both methods can be found in supplementary Sec. II. In both methods, we quote the noise assuming phase-preserving mode of amplification, with the pump frequency detuned from the measurement frequency, for which the ideal amplifier would have added noise photoncaves1982quantum.
@@
We start by using the Y-factor technique with a variable temperature source (VTS) Malnou2024low, consisting of a 50 matched microwave attenuator thermally anchored to a copper bracket with a heater and ruthenium oxide thermometer, and thermally isolated from the mixing chamber stage of the dilution refrigerator. By varying the VTS temperature, thermal noise is generated corresponding to a known photon occupation number , where , are Planck and Boltzmann constants, is the frequency and is the attenuator temperature. Fig. 3(a) shows the measured output noise power versus frequency for various . Unlike the expectation for an ideal linear amplifier where would increase with , we observe a reduction in output noise with increasing temperatures. As discussed in Ref. Malnou2024low, this behavior arises because of gain compression, from the broadband Johnson noise at elevated saturating the amplifier and reducing its gain.
@@
This gain compression is visualized in Fig. 3(b), where we plot the gain versus frequency measured by the VNA for each . The gain decreases with increasing , dropping from 21.6 dB at 0.1 K to 12.9 dB at 1.75 K. To correctly extract the added noise, we account for this gain reduction by scaling the noise spectrum with the measured gain following the procedure discussed in Ref. Malnou2024low, and further described in supplementary Sec. III. The resulting WJPA-added noise as a function of frequency is shown in Fig. 3(c). It shows that across the amplifier bandwidth, is approximately 2 photon. It is worth noting that this value likely overestimates the intrinsic JPA performance due to the lossy stainless steel cable (2.1 dB at 22 GHz, see wiring details in Supplementary Fig. S2) connecting the VTS to the WJPA. Because the reference plane for the Y-factor analysis is defined at the VTS, any attenuation from the intervening cable effectively appear as additional added noise. The finite temperature gradient across the cable length also makes the reference plane ambiguous. Considering these effects, as well as other approximations discussed in Supplementary Sec. III, we quote photons as a conservative result of the WJPA-added noise.
@@
Next, we implemented a qubit-based noise calibration approach using a transmon-cavity system macklin2015near; Bultink2018general with a readout cavity frequency of GHz. Details of this method are described in supplementary Sec. IV. Fig. 3(d) shows the calibrated noise spectrum referenced to the input of the WJPA, acquired at 22 GHz. When the WJPA is turned off, the noise power is dominated by the HEMT noise, measured to be approximately 10.3 photons, or system noise temperature K at 22 GHz. With the WJPA turned on, the total system added noise drops to photon, or system noise temperature K. Note that combines the noise contributions of losses from cables and components between the transmon-cavity and WJPA, the WJPA-added noise , and the residual added-noise of the rest of the amplification chain after the WJPA. We can also express these results in terms of a system quantum measurement efficiencyMallet2011quantum; white2023 with the WJPA turned on, compared to 4.6% with the WJPA turned off.
@@
To summarize, both noise calibration methods indicate the system added noise with the WJPA operating of approximately 2 photon. This value is likely limited by losses in coaxial cables and circulators between the VTS/qubit measurement plane and the WJPA port, which could be improved with high-frequency microwave system engineering in the future. Overall, this noise performance confirms the WJPA’s capability to operate near the quantum limit and substantially improve fidelity of measurement in high-frequency readout applications.
@@
In summary, we have demonstrated a wireless, waveguide-coupled Josephson parametric amplifier (WJPA) that features a lumped-element resonator embedded with a JJ SQUID and coupled to the input/output port via a lithographically defined waveguide-to-slotline coupling structure. Operated in four-wave mixing mode with a single pump tone, the WJPA achieves over 20 dB of gain across a tunable range from 21 GHz to 23.5 GHz, with typical 3 MHz bandwidth. Using both Y-factor measurements and a qubit-based noise calibration, we extract a system added noise of approximately two photons. These results represent the first demonstration of a Josephson parametric amplifier operating above 20 GHz.
Our work opens up a few natural next steps. The measured P dBm at the operating frequency of 22.9 GHz can be alternatively expressed as 21 photons/s. This number should be compared to typical drive power values used for fast, high-fidelity qubit readout photons/sHeinsoo2018; Hazra2025; Spring2025. If as expected, similar drive powers are also needed for high frequency qubit readout, then amplifier compression could pose a limit to readout fidelity. Thus a natural next step to our work would be to characterize the fidelity and repeatability of high-frequency dispersive qubit readout with such an amplifier. We also note that strategies to increase amplifier compression power by arraying junctions and increasing their critical currentEichler2014; Frattini2018; Kaufman2025Simple, as well as increasing bandwidth through impedance engineeringNaaman2022, should be adaptable to higher operating frequency as well, as long as attention is paid to the limits imposed by the junction plasma frequency.
Furthermore, while this work realized a WJPA with Al/AlOx junctions, the design is readily implementable with Josephson elements realized with superconductors such as niobium that can operate at high temperature Kanferov2024millimeter, as well as kinetic inductance materials like niobium nitrideAnferov2020; xu2023magnetic, NbTiNParker2022deg; hung2025broadband, and granular aluminumMaleeva2018GrAl; Zapata2024granular. In particular, we note that self-resonance frequencies in excess of 70 GHz have been measured with granular aluminumMaleeva2018GrAl. Characterizing such self-resonance frequencies for other Josephson elements and kinetic inductance materials will be especially helpful in understanding the limits on designing amplifiers for higher operation frequencies.